OpenStudy (anonymous):

please help Which sequence could be partially defined by the recursive formula f (n + 1) = f(n) + 2.5 for n ≥ 1? 2.5, 6.25, 15.625, 39.0625, … 2.5, 5, 10, 20 –10, –7.5, –5, –2.5, … –10, –25, 62.5, 156.25

OpenStudy (danjs):

each next term is 2.5 more than the last term

OpenStudy (anonymous):

So would the answer be the last one? @DanJS

OpenStudy (danjs):

sorry i had to run for a few min

OpenStudy (danjs):

the next term is the term before +2.5 -10 + 2.5 = -7.5 -7.5 + 2.5 = -5 -5 + 2.5 = -2.5